# Elusive groups from non-split extensions

### v1

@@ Abstract.
A finite transitive permutation group is elusive if it contains no derangements of prime order. These groups are closely related to a longstanding open problem in algebraic graph theory known as the Polycirculant Conjecture, which asserts that no elusive group is $2$-closed. Existing constructions of elusive groups mostly arise from split extensions. In this paper, we initiate the construction of elusive groups via non-split extensions. This approach offers a key advantage: it allows the construction of elusive groups even when the acting group is not elusive. As a demonstration, we construct elusive groups of new degrees, namely $p^{3k-4}(p+1)/2$ for each Mersenne prime $p\geq 7$ and integer $k\geq 2$, along with the first odd degree $225$ and the first twice odd degree $450$. We conclude by proposing further problems to advance this new direction of research.
@@ 1. Introduction
$p^{k}2^{n}q_{1}\cdots q_{r}\ \text{ and }\ 7^{k}2^{2}3,$ (1)
@@
There exists an elusive group of degree $5^{2}3^{2}$ which has the form $(\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ and stabilizer $(\mathrm{C}_{3}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}^{2}$, and there exists an elusive group of degree $5^{2}3^{2}2$ which has the form $(\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ and stabilizer $(\mathrm{C}_{3}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}$.
@@ 2. Illustration of idea
Let $a$ be an element (if any) of order $p$ in $X$. Since $|N|$ is not divisible by $p$, the order of $\overline{a}$ is also $p$. Since $G$ is $p$-elusive with point stabilizer $H$, there exists $g\in G$ such that $\overline{a}^{g}\in H$. This implies the existence of $b\in X$ such that $\overline{a^{b}}\in H=\overline{Y}$, that is, $a^{b}\in NY$. Again since $|N|$ is not divisible by $p$, every Sylow $p$-subgroup $P$ of $Y$ is a Sylow $p$-subgroup of $NY$. Hence there exists $c\in NY$ such that $a^{bc}\in P\leq Y$. This shows that $X$ is $p$-elusive. ∎
By Lemma 2.1, to construct elusive groups as group extensions $N.G$ from a transitive permutation group $G$, we need only to ensure that $N.G$ is $p$-elusive for every prime $p$ satisfying one of the following:
@@
It turns out that, in the setting of Lemma 2.3, it is still possible that $N.G$ is elusive when $|N|$ is divisible by $p$. For a prime $p$ and transitive permutation groups $G$ and $N.G$, we say that $N.G$ overcomes $p$ for $G$ if $N.G$ is $p$-elusive while $G$ is not. Let us illustrate the possibility of such a situation by proving the following special case of Theorem 1.3.
@@
Let $G\cong\mathrm{PSL}_{2}(p)$ for some Mersenne prime $p\geq 7$, and let $H\cong\mathrm{D}_{p-1}$ be a subgroup of $G$, as in Example 2.2. Then $G$ is $p^{\prime}$-elusive with stabilizer $H$. Consider the action of $G$ on $N:=\mathrm{C}_{p}^{3}$ by viewing $G=\Omega_{3}(p)$ acting on the space $N=\mathbb{F}_{p}^{3}$. Then $H$ is $\Omega_{2}^{+}(p){:}\mathrm{C}_{2}$ and stabilizes a $2$-dimensional subspace $M$ of $N$. Let $X=N.G=\mathrm{C}_{p}^{3}.\mathrm{PSL}_{2}(p)$ be an extension of $N$ by $G$ with this action of $G$, and let $Y=M.H=M{:}H=\mathrm{C}_{p}^{2}{:}\mathrm{D}_{p-1}$ be a subgroup of $\mathbf{N}_{X}(M)=N.H=N{:}H$ (note that $|N|$ and $|H|$ are coprime).
@@
The group (3) is III(xi) in the table of groups of order $p^{n}$ in Burnside’s 1897 book [6, §73]. In particular, it has order $p^{4}$, and so does $W$.
@@
Let $W$ be the group defined in (2) with odd prime $p$. Then $W\setminus\langle a,b,c\rangle$ has no elements of order $p$.
@@
$d^{p}(s^{i})^{p}=(ds^{i})^{p}c_{2}^{\binom{p}{2}}\dots c_{p-1}^{\binom{p}{p-1}}c_{p}$
@@ 3. Elusive groups of degree $p^{3k-4}(p+1)/2$
if $a_{i,j}\equiv b_{i,j}\ (\bmod\,p^{\ell})$ for each $i,j\in\{1,2\}$.
@@ 4. Elusive groups of degree not divisible by 44
Its transitive representation of degree $15$ with stabilizer a Sylow 2-subgroup $C_{2}^{2}$ is $2$-elusive, but neither $3$- nor $5$-elusive. Let $U=\mathbb{F}_{3}^{4}$ and $V=\mathbb{F}_{5}^{3}$ be the fully deleted permutation modules of $\mathrm{A}_{5}$ over $\mathbb{F}_{3}$ and $\mathbb{F}_{5}$, respectively, which are both (absolutely) irreducible modules of $\mathrm{A}_{5}$ [23, Page 186]. According to [24, Corollary 1] and [25], both $\mathrm{H}^{2}(\mathrm{A}_{5},U)$ and $\mathrm{H}^{2}(\mathrm{A}_{5},V)$ have dimension $1$. Hence, it may be possible to construct an extension of $\mathrm{A}_{5}$ that overcomes $3$ and $5$ for $\mathrm{A}_{5}$. We do so in the rest of this section, leading to a proof of Theorem 1.4.
With $\bigoplus_{i=1}^{5}\mathbb{F}_{3}e_{i}$ being the permutation module of $\mathrm{A}_{5}$ over $\mathbb{F}_{3}$, take
@@
Similarly, if $\bigoplus_{i=1}^{5}\mathbb{F}_{5}e_{i}$ is the permutation module of $\mathrm{A}_{5}$ over $\mathbb{F}_{5}$, then taking
@@
$V=\bigoplus_{i=1}^{3}\mathbb{F}_{5}v_{i}$ with
@@
Inspired by (7)–(11), we construct as follows a group $X$ as an extension of $U\times V$ that can overcome $3$ and $5$ for $\mathrm{A}_{5}$.
Construction 4.1.
Let $X$ be the group with generators $u_{1},u_{2},u_{3},u_{4},v_{1},v_{2},v_{3},x,y$ and defining relations $u_{i}^{3}=[u_{i},u_{j}]=v_{i}^{5}=[v_{i},v_{j}]=[u_{i},v_{j}]=1$ for all admissible $i$ and $j$,
@@
and those in (8), (9), (10), (11). Then $U:=\langle u_{1},u_{2},u_{3},u_{4}\rangle=\mathrm{C}_{3}^{4}$ and $V:=\langle v_{1},v_{2},v_{3}\rangle=\mathrm{C}_{5}^{3}$ are both normal subgroups of $X$. Let $N=UV=U\times V$, let $\overline{\phantom{\varphi}}\,$ be the quotient map from $X$ to $G:=X/N$, and let $M=\langle u_{1}u_{2},u_{2}u_{3}^{-1},u_{3}u_{4},v_{1}v_{2},v_{2}v_{3}^{-1}\rangle=\mathrm{C}_{3}^{3}\times\mathrm{C}_{5}^{2}$. Since it is straightforward to verify that $\langle\overline{x},\overline{y}[\overline{x},\overline{y}]^{2}\rangle=\mathrm{C}_{2}^{2}$ is a subgroup of $G$ stabilizing $M$, there exists a subgroup $Y=M{:}H=(\mathrm{C}_{3}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}^{2}$ of $X$ with $H\cong\overline{H}=\langle\overline{x},\overline{y}[\overline{x},\overline{y}]^{2}\rangle$. ∎
@@
Although not needed in the sequel, it can be directly verified that one may take $H=\langle x,u_{1}^{-1}u_{2}^{-1}u_{3}u_{4}v_{1}^{-1}v_{2}^{-2}v_{3}^{2}y[x,y]^{2}\rangle$ in Construction 4.1.
The group $G=X/N=\langle\overline{x},\overline{y}\rangle$ in Construction 4.1 satisfies $\overline{x}^{2}=\overline{y}^{3}=(\overline{x}\overline{y})^{5}=1$ and hence is isomorphic to $\mathrm{A}_{5}$ by (7). In fact, $\overline{x}\mapsto(1,2)(3,4)$ and $\overline{y}\mapsto(1,3,5)$ give an isomorphism from $G$ to $\mathrm{A}_{5}$. It follows that $X=N.G=(\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}).\mathrm{A}_{5}$. We will see in the next lemma that $X$ is a non-split extension of $N=\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}$ by $G=\mathrm{A}_{5}$. Note from (8)–(11) that $U$ and $V$ are fully deleted permutation modules of $G$ and so are irreducible. Hence $U$ and $V$ are minimal normal in $X$.
@@
With the notation in Construction 4.1, $Y$ is core-free in $X=(\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$, and $X\setminus N$ has no elements of order $3$ or $5$.
@@
Since $U$ and $V$ are minimal normal subgroups of $X$ while $M$ contains neither $U$ nor $V$, the subgroup $Y=M{:}\mathrm{C}_{2}^{2}$ is core-free in $X$. Consider the Sylow $3$-subgroup $P=U\langle y\rangle$ and Sylow $5$-subgroup $Q=V\langle xy\rangle$ of $X$. As in the proof of Lemma 2.7, we see that there are no elements of order $3$ in $P\setminus U$ and no elements of order $5$ in $Q\setminus V$. Thus, there are no elements of order $3$ or $5$ in $X\setminus N$. In particular, $X=(\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$, which completes the proof. ∎
@@
$M\cap U=\langle u_{1}u_{2},u_{2}u_{3}^{-1},u_{3}u_{4}\rangle=\langle u_{1}u_{2},u_{2}u_{4}\rangle\times\langle u_{1}u_{2}^{-1}u_{3}^{-1}\rangle,$ (12)

### v2

@@ Abstract.
A finite transitive permutation group is elusive if it contains no derangements of prime order. These groups are closely related to a longstanding open problem in algebraic graph theory known as the Polycirculant Conjecture, which asserts that no elusive group is $2$-closed. Existing constructions of elusive groups mostly arise from split extensions. In this paper, we initiate the construction of elusive groups via non-split extensions. As a demonstration, we construct elusive groups of new degrees, namely $p^{3k-4}(p+1)/2$ for each Mersenne prime $p\geq 7$ and integer $k\geq 2$. We also construct the first examples of elusive groups with odd degree, namely $3^{k+1}\cdot 5^{2}$, and twice odd degree, namely $2\cdot 3^{k+1}\cdot 5^{2}$ for each $k\geq 1$. We conclude by proposing further problems to advance this new direction of research.
@@ 1. Introduction
$2^{2}\cdot 3\cdot 7^{k}\ \text{ and }\ 2^{n}\cdot p^{k}\cdot q_{1}\cdots q_{r},$ (1)
@@
Remark.
The group $\mathrm{C}_{p}^{3}.\mathrm{C}_{p}^{3}.\cdots.\mathrm{C}_{p}^{3}$ in Theorem 1.3 is nonabelian for $k\geq 3$, as shown in the remark after the proof of Lemma 3.2.
@@
For every integer $k\geq 1$, there exists an elusive group of degree $3^{k+1}\cdot 5^{2}$ that has the form $(\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ and stabilizer $(\mathrm{C}_{3^{k}}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}^{2}$, and there exists an elusive group of degree $2\cdot 3^{k+1}\cdot 5^{2}$ that has the form $(\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ and stabilizer $(\mathrm{C}_{3^{k}}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}$.
@@ 2. Illustration of idea
Let $a$ be an element (if any) of order $p$ in $X$. Since $|N|$ is not divisible by $p$, the order of $\overline{a}$ is also $p$. Since $G$ is $p$-elusive with point stabilizer $H$, there exists $g\in G$ such that $\overline{a}^{g}\in H$. This implies the existence of $b\in X$ such that $\overline{a^{b}}\in H=\overline{Y}$, that is, $a^{b}\in NY$. Again, since $|N|$ is not divisible by $p$, every Sylow $p$-subgroup $P$ of $Y$ is a Sylow $p$-subgroup of $NY$. Hence there exists $c\in NY$ such that $a^{bc}\in P\leq Y$. This shows that $X$ is $p$-elusive. ∎
By Lemma 2.1, to construct elusive groups as group extensions $N.G$ from a transitive permutation group $G$, we need only ensure that $N.G$ is $p$-elusive for every prime $p$ satisfying one of the following:
@@
In the setting of Lemma 2.3, it is still possible that $N.G$ is elusive when $|N|$ is divisible by $p$. For a prime $p$ and transitive permutation groups $G$ and $N.G$, we say that $N.G$ overcomes $p$ for $G$ if $N.G$ is $p$-elusive while $G$ is not. Let us illustrate the possibility of such a situation by proving the following special case of Theorem 1.3.
@@
Let $G\cong\mathrm{PSL}_{2}(p)$ for some Mersenne prime $p\geq 7$, and let $H\cong\mathrm{D}_{p-1}$ be a subgroup of $G$, as in Example 2.2. Then $G$ is $p^{\prime}$-elusive with stabilizer $H$. Consider the action of $G$ on $N\coloneqq\mathrm{C}_{p}^{3}$ by viewing $G=\Omega_{3}(p)$ acting on the space $N=\mathbb{F}_{p}^{3}$. Then $H$ is $\Omega_{2}^{+}(p){:}\mathrm{C}_{2}$ and stabilizes a $2$-dimensional subspace $M$ of $N$. Let $X=N.G=\mathrm{C}_{p}^{3}.\mathrm{PSL}_{2}(p)$ be an extension of $N$ by $G$ with this action of $G$, and let $Y=M.H=M{:}H=\mathrm{C}_{p}^{2}{:}\mathrm{D}_{p-1}$ be a subgroup of $\mathbf{N}_{X}(M)=N.H=N{:}H$ (note that $|N|$ and $|H|$ are coprime).
@@
The group (3) is III(xi) in the table of groups of order $p^{n}$ in [6, §73]. In particular, it has order $p^{4}$, and so does $W$.
@@
Let $W$ be the group defined in (2) for an odd prime $p$. Then $W\setminus\langle a,b,c\rangle$ has no elements of order $p$.
@@
$d^{p}(s^{i})^{p}=(ds^{i})^{p}c_{2}^{\binom{p}{2}}\cdots c_{p-1}^{\binom{p}{p-1}}c_{p}$
@@ 3. Elusive groups of degree $p^{3k-4}(p+1)/2$
if $a_{i,j}\equiv b_{i,j}\ (\bmod\,p^{\ell})$ for each $i,j\in\{1,2\}$. While the following lemma is closely related to results in [20, III§17], we provide a self-contained treatment here to establish the specific properties required for our analysis.
@@
Remark.
For $k\geq 3$, the group $\widehat{N}$ is nonabelian. This can be seen, for example, from the following calculation:
$\begin{pmatrix}1+p&p\\ -p&1-p\end{pmatrix}\begin{pmatrix}1&p\\ 0&1\end{pmatrix}$ $=\begin{pmatrix}1+p&2p+p^{2}\\ -p&1-p-p^{2}\end{pmatrix}$
$\begin{pmatrix}1&p\\ 0&1\end{pmatrix}\begin{pmatrix}1+p&p\\ -p&1-p\end{pmatrix}$ $=\begin{pmatrix}1+p-p^{2}&2p-p^{2}\\ -p&1-p\end{pmatrix}$.
@@ 4. Elusive groups of degree not divisible by 44
Its transitive representation of degree $15$ with stabilizer a Sylow 2-subgroup $\mathrm{C}_{2}^{2}$ is $2$-elusive, but neither $3$- nor $5$-elusive. Let $U=\mathbb{F}_{3}^{4}$ and $V=\mathbb{F}_{5}^{3}$ be the fully deleted permutation modules of $\mathrm{A}_{5}$ over $\mathbb{F}_{3}$ and $\mathbb{F}_{5}$, respectively, which are both (absolutely) irreducible modules of $\mathrm{A}_{5}$ [23, p. 186]. According to [24, Corollary 1] and [25], both $\mathrm{H}^{2}(\mathrm{A}_{5},U)$ and $\mathrm{H}^{2}(\mathrm{A}_{5},V)$ have dimension $1$. Hence, it may be possible to construct an extension of $\mathrm{A}_{5}$ that overcomes $3$ and $5$ for $\mathrm{A}_{5}$. In this section, we construct infinitely many such extensions, leading to a proof of Theorem 1.4.
If $\bigoplus_{i=1}^{5}\mathbb{F}_{3}e_{i}$ is the permutation module of $\mathrm{A}_{5}$ over $\mathbb{F}_{3}$, then take
@@
Similarly, if $\bigoplus_{i=1}^{5}\mathbb{F}_{5}e_{i}$ is the permutation module of $\mathrm{A}_{5}$ over $\mathbb{F}_{5}$, then take
@@
so $V=\bigoplus_{i=1}^{3}\mathbb{F}_{5}v_{i}$ with
@@
Inspired by (7)–(11), we construct as follows a group $X$ as an extension of $U\times V$ that can overcome $3$ and $5$ for $\mathrm{A}_{5}$. Fix an integer $k\geq 1$. Consider the group $X$ with generators $u_{1}$, $u_{2}$, $u_{3}$, $u_{4}$, $v_{1}$, $v_{2}$, $v_{3}$, $x$, $y$ and defining relations $u_{i}^{3^{k}}=[u_{i},u_{j}]=v_{i}^{5}=[v_{i},v_{j}]=[u_{i},v_{j}]=1$ for all admissible $i$ and $j$,
@@
and those in (8), (9), (10), (11). Let $U=\langle u_{1},u_{2},u_{3},u_{4}\rangle$ and $V=\langle v_{1},v_{2},v_{3}\rangle$ and $N=\langle U,V\rangle$.
Lemma 4.1.
For every integer $k\geq 1$, the group $X$ has the form $(\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}).\mathrm{A}_{5}$.
Proof.
Based on the relations of $X$ satisfied by the generators $u_{1},u_{2},u_{3},u_{4}$ and $v_{1},v_{2},v_{3}$, it is clear that $U$ and $V$ are quotients of $\mathrm{C}_{3^{k}}^{4}$ and $\mathrm{C}_{5}^{3}$, respectively. In particular, $U$ and $V$ must intersect trivially. Moreover, both $U$ and $V$ are normal subgroups of $X$. It follows that $N=\langle U,V\rangle=UV=U\times V$. Since $x^{2},y^{3},(xy)^{5}\in N$, the group $X/N=X/(U\times V)$ is isomorphic to $\mathrm{A}_{5}$ by (7). Hence, $X\cong(U\times V).\mathrm{A}_{5}$. Therefore, in order to prove the lemma, it suffices to show that $|U|=3^{4k}$ and $|V|=5^{3}$.
Let $X_{U}$ and $X_{V}$ denote the quotients of $X$ by $U$ and $V$, respectively, each admitting a presentation obtained from the presentation of $X$ by setting the generators $u_{1},u_{2},u_{3},u_{4}$ and $v_{1},v_{2},v_{3}$ to $1$, respectively. One can directly check (for example, by Magma [4]) that $X_{U}$ is isomorphic to ${\mathrm{C}_{5}^{3}}.\mathrm{A}_{5}$. As a consequence, $|V|=5^{3}$. To complete the proof, we show that $X_{V}$ is isomorphic to ${\mathrm{C}_{3^{k}}^{4}}.\mathrm{A}_{5}$. We do this by specifying an embedding $\phi$ of $X_{V}=\langle u_{1},u_{2},u_{3},u_{4},x,y\rangle$ into the wreath product $W\coloneqq\mathrm{C}_{3^{k+1}}\wr\mathrm{A}_{5}=\mathrm{C}_{3^{k+1}}^{5}\rtimes\mathrm{A}_{5}$.
Let $b_{1},\ldots,b_{5}$ be the natural generators of the base group $\mathrm{C}_{3^{k+1}}^{5}$ of $W$. For $i\in\{1,2,3,4\}$, let
$h_{i}=b_{i}^{3}b_{5}^{-3}.$
Let $r=(1,2)(3,4)$ and $s=(1,3,5)$ be elements of $A_{5}$, and let
$\rho=b_{3}^{-1}b_{4}r\ \ \text{and}\ \ \sigma=b_{2}b_{4}^{-1}s$
be elements of $W$. We first verify that $\phi\colon X_{V}\rightarrow W$ given by $\phi(u_{i})=h_{i}$ for $i\in\{1,2,3,4\}$, $\phi(x)=\rho$ and $\phi(y)=\sigma$ is a group homomorphism.
Obviously, $h_{i}^{3^{k}}=1$ and $h_{i},h_{j}]=1$ for all $i,j\in\{1,2,3,4\}$. Observe that
$\rho^{2}$ $=b_{3}^{-1}b_{4}(b_{3}^{-1}b_{4})^{r}r^{2}=b_{3}^{-1}b_{4}b_{4}^{-1}b_{3}=1,$
$\sigma^{3}$ $=(b_{2}b_{4}^{-1})^{3}s^{3}=b_{2}^{3}b_{4}^{-3}=b_{2}^{3}b_{5}^{-3}(b_{4}^{3}b_{5}^{-3})^{-1}=h_{2}h_{4}^{-1},$
so the relations $\rho^{2}=1$ and $\sigma^{3}=h_{2}h_{4}^{-1}$ hold. To verify the relation $(\rho\sigma)^{5}=1$, we compute
$\rho\sigma=b_{3}^{-1}b_{4}rb_{2}b_{4}^{-1}s$ $=b_{3}^{-1}b_{4}(b_{2}b_{4}^{-1})^{r}rs$
$=b_{3}^{-1}b_{4}b_{1}b_{3}^{-1}(1,2)(3,4)(1,3,5)=b_{1}b_{3}^{-2}b_{4}(1,2,3,4,5)$
and write $\rho\sigma=bt$ with $b=b_{1}b_{3}^{-2}b_{4}$ and $t=(1,2,3,4,5)$. Then
$(\rho\sigma)^{5}=(bt)^{5}=bb^{t^{-1}}b^{t^{-2}}b^{t^{-3}}b^{t^{-4}}t^{5}=bb^{t}b^{t^{2}}b^{t^{3}}b^{t^{4}},$
and since
$b^{t}=b_{2}b_{4}^{-2}b_{5},\ \ b^{t^{2}}=b_{3}b_{5}^{-2}b_{1},\ \ b^{t^{3}}=b_{4}b_{1}^{-2}b_{2},\ \ b^{t^{4}}=b_{5}b_{2}^{-2}b_{3},$
it follows that $(bt)^{5}=1$. In particular, $(\rho\sigma)^{5}=1$. Now we establish the compatibility of the generators $h_{1},h_{2},h_{3},h_{4},\rho,\sigma$ with relations (8) and (9). Let $i\in\{1,2,3,4\}$. Then
$h_{i}^{\rho}=\rho h_{i}\rho=b_{3}^{-1}b_{4}rb_{i}^{3}b_{5}^{-3}b_{3}^{-1}b_{4}r$ $=b_{3}^{-1}b_{4}(b_{i}^{3}b_{5}^{-3}b_{3}^{-1}b_{4})^{r}$
$=b_{3}^{-1}b_{4}b_{i^{r}}^{3}b_{5}^{-3}b_{4}^{-1}b_{3}=b_{i^{r}}^{3}b_{5}^{-3}=h_{i^{r}},$
as in (8). For the relation corresponding to (9),
$h_{i}^{\sigma}=b_{2}^{-1}b_{4}s^{-1}b_{i}^{3}b_{5}^{-3}b_{2}b_{4}^{-1}s$ $=b_{2}^{-1}b_{4}(b_{i}^{3}b_{5}^{-3}b_{2}b_{4}^{-1})^{s}$
$=b_{2}^{-1}b_{4}b_{i^{s}}^{3}b_{1}^{-3}b_{2}b_{4}^{-1}=b_{1}^{-3}b_{i^{s}}^{3}=h_{1}^{-1}b_{i^{s}}^{3}b_{5}^{-3},$
from which it follows immediately that
$h_{1}^{\sigma}=h_{1}^{-1}b_{3}^{3}b_{5}^{-3}=h_{1}^{-1}h_{3}$ $,\quad h_{2}^{\sigma}=h_{1}^{-1}b_{2}^{3}b_{5}^{-3}=h_{1}^{-1}h_{2},$
$h_{3}^{\sigma}=h_{1}^{-1}b_{5}^{3}b_{5}^{-3}=h_{1}^{-1}$ $,\quad h_{4}^{\sigma}=h_{1}^{-1}b_{4}^{3}b_{5}^{-3}=h_{1}^{-1}h_{4},$
as required.
Thus far we have verified that $\phi$ is a group homomorphism from $X_{V}$ to $W$. Let $H=\langle h_{1},h_{2},h_{3},h_{4}\rangle$ and $G=\langle H,\rho,\sigma\rangle$ be the image of $\phi$. It is clear that $H=\mathrm{C}_{3^{k}}^{4}$. Note that $H$ is normal in $G$ (in fact, $H$ is normal in $W$) with the quotient $G/H$ isomorphic to $A_{5}$ (generated by $r=(1,2)(3,4)$ and $s=(1,3,5)$). Hence, $G=\mathrm{C}_{3^{k}}^{4}.\mathrm{A}_{5}$, and so $|X_{U}|\geq|G|=3^{4k}\times 60$. As previously observed, $U$ is a quotient of $\mathrm{C}_{3^{k}}^{4}$ and in particular, $3^{4k}\geq|U|$. Finally, $U$ is normal in $X_{V}$ and $X_{V}/U\cong A_{5}$. In particular, $|X_{V}|=|U||A_{5}|\leq 3^{4k}\times 60$. It follows that $|X_{U}|=3^{4k}\times 60=|G|$ and consequently, $\phi$ is a group isomorphism. ∎
Construction 4.2.
Fix an integer $k\geq 1$. Let $X$, $U$, $V$ and $N$ be as defined above, let $\overline{\phantom{\varphi}}\,$ be the quotient map from $X$ to $G\coloneqq X/N$, and let
$M=\langle u_{1}u_{2},u_{2}u_{3}^{-1},u_{3}u_{4},v_{1}v_{2},v_{2}v_{3}^{-1}\rangle.$
@@
Note that $M\cong\mathrm{C}_{3^{k}}^{3}\times\mathrm{C}_{5}^{2}$. Since it is straightforward to verify that $\langle\overline{x},\overline{y}[\overline{x},\overline{y}]^{2}\rangle=\mathrm{C}_{2}^{2}$ is a subgroup of $G$ stabilizing $M$, there exists a subgroup $Y=M{:}H=(\mathrm{C}_{3^{k}}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}^{2}$ of $X$ with
$H\cong\overline{H}=\langle\overline{x},\overline{y}[\overline{x},\overline{y}]^{2}\rangle.$
Let $W=\langle M,x\rangle=M{:}\mathrm{C}_{2}$ be a subgroup of $Y$.
@@
Although not needed in the sequel, it can be directly verified that one may take $H=\langle x,u_{1}^{-1}u_{2}^{-1}u_{3}u_{4}v_{1}^{-1}v_{2}^{-2}v_{3}^{2}y[x,y]^{2}\rangle$ in Construction 4.2.
By Lemma 4.1, for a fixed integer $k\geq 1$, the group $X$ from Construction 4.2 satisfies $X=N.G=(\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}).\mathrm{A}_{5}$. We will see in the next lemma that $X$ is a non-split extension of $N=\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}$ by $G=\mathrm{A}_{5}$.
@@
For every integer $k\geq 1$, both $Y$ and $W$ are core-free in $X=(\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$, and $X\setminus N$ has no elements of order $3$ or $5$.
@@
[…]
Let $J\cong\mathrm{C}_{3}^{4}$ be the subgroup of $N$ generated by the elements of order $3$ in $N$. To prove that $Y$ and $W$ are core-free in $X$, it suffices to show that the minimal normal subgroups of $X$ are precisely $J$ and $V$.
From the defining relations of $X$ in Construction 4.2, it is easy to see that the action of $X$ on $J$ and $V$ corresponds to the action of $\mathrm{A}_{5}$ on its fully deleted permutation modules over $\mathbb{F}_{3}$ and $\mathbb{F}_{5}$, respectively, which are irreducible. Therefore, both $J$ and $V$ are minimal normal subgroups of $X$. We next show that they are the only ones.
Let $K$ be a minimal normal subgroup of $X$. Since the quotient of $X$ by $N$ is a simple group, either $X=NK$ or $K\leq N$. If $X=NK$, then the minimality of $K$ implies $N\cap K=1$ and so X=N:KX=N{:}K, contradicting the conclusion that $X$ is a non-split extension of $N$ by $\mathrm{A}_{5}$. Thus, $K\leq N$. Note that a nontrivial element of $N$ is in $J$ if and only if it has order $3$, and is in $V$ if and only if it has order $5$. Therefore, $K$ must intersect $J$ or $V$ nontrivially. Hence $K=J$ or $V$, as desired.
[…]
Suppose for a contradiction that $uy^{i}\in P$ has order $3$ for some $u\in U$ and $i\in\{1,2\}$. Again by the Hall-Petrescu identity [20, Theorem III.9.4], we obtain
$u^{3}(y^{i})^{3}=(uy^{i})^{3}c_{2}^{3}c_{3}=c_{2}^{3}c_{3}$
for some element $c_{j}$ in the $j$th term $P_{j}$ of the lower central series of $P$, where $j\in\{2,3\}$. Since $P_{3}\leq P_{2}=P^{\prime}$, it follows that
$u^{3}(y^{i})^{3}\in P^{\prime}\leq\langle u_{1},u_{3}\rangle.$ (12)
Expressing $u=u_{1}^{a}u_{2}^{b}u_{3}^{c}u_{4}^{d}$ with $a,b,c,d\in\mathbb{Z}$ in (12) and recalling that $y^{3}=u_{2}u_{4}^{-1}$, we derive
$u_{1}^{3a}u_{2}^{3b+i}u_{3}^{3c}u_{4}^{3d-i}\in\langle u_{1},u_{3}\rangle.$
This implies that $u_{2}^{3b+i}u_{4}^{3d-i}\in\langle u_{1},u_{3}\rangle$, which is not possible as $i\in\{1,2\}$. The proof is thus complete. ∎
@@
$M\cap J=\langle w_{1}w_{2},w_{2}w_{3}^{-1},w_{3}w_{4}\rangle=\langle w_{1}w_{2},w_{2}w_{4}\rangle\times\langle w_{1}w_{2}^{-1}w_{3}^{-1}\rangle,$ (13)